# Contrastive Learning and Mixture of Experts Enables Precise Vector Embeddings

### v2

@@ Abstract
The advancement of transformer neural networks has significantly elevated the capabilities of sentence similarity models, but they struggle with highly discriminative tasks and produce sub-optimal representations of important documents like scientific literature. With the increased reliance on retrieval augmentation and search, representing diverse documents as concise and descriptive vectors is crucial. This paper improves upon the vectors embeddings of scientific literature by assembling niche datasets using co-citations as a similarity metric, focusing on biomedical domains. We apply a novel Mixture of Experts (MoE) extension pipeline to pretrained BERT models, where every multi-layer perceptron section is enlarged and copied into multiple distinct experts. Our MoE variants perform well over $N$ scientific domains with $N$ dedicated experts, whereas standard BERT models excel in only one domain. Notably, extending just a single transformer block to MoE captures 85% of the benefit seen from full MoE extension at every layer. This holds promise for versatile and efficient One-Size-Fits-All transformer networks for numerically representing diverse inputs. Our methodology marks significant advancements in representing scientific text and holds promise for enhancing vector database search and compilation.
@@ 1 Introduction
Fortunately, several advancements have paved the way toward effective sentence similarity models over an arbitrary number of domains. Work from the metascience community has introduced co-citation networks as an easy way to gather many papers with a high degree of similarity [11, 12, 13, 14, 15, 16, 17, 18, 19]. While deep learning provides Mixture of Experts (MoE) models which have a promising multidomain / multitask learning architecture [20]. Taking advantage of these methods, we propose the following framework to build discriminative vector representations of scientific papers from abstracts alone:
@@ 2.1 Data Compilation
Our validation dataset was constructed from the remaining training dataset. The remaining training dataset was split randomly in a 99:1 ratio without duplicates, with the larger of the new datasets being used as our final training dataset and the smaller of the two being used as the initial validation dataset. On top of this initial validation dataset of similar abstract pairs, an equal amount of dissimilar abstract pairs were added. We generated pairs of dissimilar papers by compiling pairs of papers cited at least 15 times each but had never been co-cited together. While it is impossible to guarantee any two non-co-cited papers will not be cited together in the future we minimized this possibility by requiring papers to be cited individually at least 15 times. Note that while we produced pairs of dissimilar papers for model evaluation, the production of dissimilar paper pairs is not necessary for model generation.
@@ 2.2.1 Mixture of Experts
Mixture of Experts (MoE) models add a linear layer or router network to each transformer block, which outputs logits from $H^{(i)}$. These logits route $H^{(i)}$ to multiple unique copies of the MLP section in the transformer block called experts [35]. For causal GPT-like models, this is typically done on a per-token basis, allowing for experts to specify in language classes like punctuation, nouns, numbers, etc [36]. For BERT-like models, this provides for sentence-wise routing of the entire $H^{(i)}$ [37]. While there are many ways to route MoE networks, two main approaches involve calling one expert per block or the top $k$ experts per block. Our variants called one expert per block, otherwise known as the switch transformer [38].
@@
For faster fine-tuning, we utilized pretrained models for this novel MoE extension approach. Our MoE extension took the MLP sections of pretrained transformers and copied them into $E$ experts with randomly initialized routers in each transformer block. In this process, an additional linear layer $W_{3}\in\mathbb{R}^{d\times I}$ and bias $b_{3}$ was added with element-wise multiplication $\odot$ to the MLP as seen in Mixtral [36], which we reasoned to be more expressive:
@@
Enforced routing was chosen for the dual CVD/COPD experiments as a proof of concept (we also did not add the additional linear layer). However, for the experiments over all five biomedical domains, we implemented the mutual information loss suggested in [39] to further leverage overlapping similarities across the gathered biomedical domains. This way, we could correlate expert activation with certain domains without direct enforcement and concatenate the top-2 expert results at each layer. Additional local experiments showed that token-wise routing results in slightly higher-end performance, even on sentence-level tasks. Thus, we use the top-2 experts per token for our final five-domain model. Because our MoE extension can be costly in terms of VRAM, we also tried an MoE extension approach with a single transformer block in the middle instead of extending all 12 - hypothesizing that much of the multidomain benefit could be achieved for a small amount of extended MoE layers.
@@ 3 Results
The performance of our models on the validation and test datasets for CVD (Table 1) and COPD (Table 2) were summarized and compared to other leading sentence similarity models, as well as a standard term frequency-inverse document frequency model (TF-IDF). We also evaluated the performance of GPT-3.5 on our generated validation and test datasets. GPT-3.5 represented an example of an LLM with a narrow representation of similar (Figure 2). GPT-3.5 regarded almost all examples that were input as dissimilar when asked whether the input examples were similar due to (correctly) identified minor differences between each input text. Given this, GPT-3.5 performance was excluded from model comparisons. We also do not show results from the MoE version of Mini as they did not score better than random on either validation dataset (0.67 $F1_{max}$).
@@ 4 Discussion
Our work advances the use of transformer language models by focusing on improving their domain-specific understanding and document-wide comprehension. We have shown that common pretrained models, including ChatGPT, cannot distinguish the differences in highly discriminative text inputs. Presumably, this phenotype in GPT-like LLMs is primarily caused by standard prompt construction and formatting. By including multiple distinct documents in a prompt, the semantic similarity of similar tokens will prevent effective distinction between the documents, as portions of each document will attend highly to each other even if they are “different” as defined by a desired discrimination. With pretrained BERT-like model usage, presumably this poor performance comes from out-of-distribution tasks and a low ability to generalize with high discrimination. The sentence similarity approach offers a more effective summarization technique by inputting documents separately and allowing the model to construct descriptive numerical representations through contrastive objectives. As vector databases become increasingly prevalent for search and retrieval, the quality of these numerical representations becomes increasingly important. Our innovative framework, which incorporates contrastive learning, custom MNR and router losses, novel special tokens, MoE seeding, and extension, as well as various routing techniques, significantly enhances sentence similarity metrics compared to pretrained transformers. We leveraged co-citation networks to construct large datasets of similar abstracts and applied our framework to scientific literature, creating nuanced representations for discriminative subfields of biomedical domains.
@@
Without applying a threshold search for the F1 scores, all pretrained models tested perform with random chance or worse when tasked with classifying papers as co-cited or not based on cosine similarity. By searching all possible cosine similarity thresholds for $F1_{max}$, pretrained models get a boost for fair comparison while creating a scenario that is unrealistic for actual inference, as a consistent similarity threshold is needed for actual use. Once compared fairly, our fine-tuned models performed better than their original counterparts. More specifically, our SciBERT SE and MoE variants performed equally on the CVD validation set with an $F1_{max}$ of 0.97, an average distance of 0.48, and an accuracy of 0.97. The Mini SE variant performed similarly with a $F1_{max}$ of 0.94, an average distance of 0.23, and an accuracy of 0.93. On the COPD validation set, Mini SE performed best with a $F1_{max}$ of 0.83, an average distance of 0.33, and an accuracy of 0.82. Following that is SciBERT SE with $F1_{max}$ 0.81 and SciBERT MoE with $F1_{max}$ 0.80. Both SciBERT variants resulted in an average distance of 0.49 and accuracies of 0.80 and 0.79, respectively. Importantly, our SciBERT variants consistently performed optimally with a high cosine similarity, implying that a standardized threshold near 0.98 or 0.99 could be utilized during inference. Further penalization terms on the average similarity between batches during training could enforce this threshold lower for different applications.
@@
Importantly, our MoE approach across all domains performs similarly to our individual-domain SE models. Additionally, the MoE seeding is fully scalable, enabling $N$ experts and $N$ new special tokens given $N$ different datasets. This is further supported by our experiments shown in Table 3, where MoE models with five experts perform well on five domains, even with a single MoE layer. The substantial improvement in performance achieved by adding a single MoE layer highlights a remarkable benefit with little added computational cost. This is particularly promising for using our single-layer MoE extension with large pretrained models to train for multitask / multidomain tasks. Future experiments may find the optimal ratio of experts per domain alongside the correct discrimination of “domain” to create a one-size-fits-all vector embedding model at the scale of Semantic Scholar.
@@
Links to all training data, model weights, and code can be found at
github.com/Gleghorn-Lab/Mixture-of-Experts-Sentence-Similarity.
@@
The authors thank Katherine M. Nelson, Ph.D., for reviewing and commenting on drafts of the manuscript. This work was partly supported by grants from the University of Delaware Graduate College through the Unidel Distinguished Graduate Scholar Award (LH), the National Institutes of Health R01HL133163 (JPG), R01HL145147 (JPG), R01DK132090 (BBK), and Indiana University (BBK).

### v3

@@ Abstract
The advancement of transformer neural networks has significantly elevated the capabilities of sentence similarity models, but they still struggle with highly discriminative tasks and may produce sub-optimal representations of important documents like scientific literature. With the increased reliance on retrieval augmentation and search, representing diverse documents as concise and descriptive vectors is crucial. This paper improves upon the vectors embeddings of scientific text by assembling niche datasets using co-citations as a similarity metric, focusing on biomedical domains. We apply a novel Mixture of Experts (MoE) extension pipeline to pretrained BERT models, where every multi-layer perceptron section is enlarged and copied into multiple distinct experts. Our MoE variants perform well over $N$ scientific domains with $N$ dedicated experts, whereas standard BERT models excel in only one domain at a time. Notably, extending just a single transformer block to MoE captures 85% of the benefit seen from full MoE extension at every layer. This holds promise for versatile and efficient One-Size-Fits-All transformer networks for numerically representing diverse inputs. Our methodology marks advancements in representation learning and holds promise for enhancing vector database search and compilation.
keywords:
Mixture of Experts | Biomedical literature | Contrastive learning | Sentence similarity
@@ 1 Introduction
Fortunately, several advancements have paved the way toward effective sentence similarity models over an arbitrary number of domains. Work from the metascience community has introduced co-citation networks as an easy way to gather many similar papers [11, 12, 13, 14, 15, 16, 17, 18, 19]. While this degree of similarity may not be perfect, co-citations have been shown to imply a high degree of similarity between papers [18]. Another promising advancement comes from the deep learning community with Mixture of Experts (MoE) models. Their learned input-dependent routing of information is a promising multidomain / multitask learning architecture without significant added overhead [20]. Taking advantage of these methods, we propose the following framework to build discriminative vector representations of scientific papers from abstracts alone:
@@ 2.1 Data Compilation
Our validation dataset was constructed from the remaining training dataset. The remaining training dataset was split randomly in a 99:1 ratio without duplicates, with the larger of the new datasets being used as our final training dataset and the smaller of the two being used as the initial validation dataset. On top of this initial validation dataset of similar abstract pairs, an equal amount of dissimilar abstract pairs were added. We generated pairs of dissimilar papers by compiling pairs of papers that had never been co-cited together. While it is impossible to guarantee any two non-co-cited papers will not be cited together in the future we minimized this possibility by requiring papers to be cited individually at least 15 times. While we produced pairs of dissimilar papers for model evaluation, the production of dissimilar paper pairs is not necessary for model training, which we discuss in our objective formulation below.
@@ 2.2.1 Mixture of Experts
Mixture of Experts (MoE) models add a linear layer or router network to each transformer block, which outputs logits from $H^{(i)}$. These logits route $H^{(i)}$ to multiple equivalent copies of the MLP section with different weights called experts [35]. In many transformer variants, this routing is typically done on a per-token basis, allowing for experts to specify in language classes like punctuation, nouns, numbers, etc [36]. We chose sentence-wise routing of the entire $H^{(i)}$ so that we could purposely structure our experts for specific domains [37]. While there are many ways to route MoE networks, two main approaches involve calling one expert per block or the top $k$ experts per block. We experiment with both approaches.
@@
For faster fine-tuning, we utilized pretrained models for this novel MoE extension approach. Our MoE extension took the MLP sections of pretrained transformers and copied them into $E$ experts with randomly initialized routers in each transformer block. In this process, an additional linear layer $W_{3}\in\mathbb{R}^{d\times I}$ and bias $b_{3}$ was added with element-wise multiplication $\odot$ to the MLP (SwiGLU activation) which has been shown to perform better than vanilla activation functions [36].
@@
Enforced routing refers to manual indexing of chosen experts, which we found worked just as well as an additional cross-entropy loss on the router logits. We chose enforced routing for the dual CVD / COPD experiments as a proof of concept (we also did not add the additional router linear layer). However, for the experiments over all five biomedical domains, we implemented the mutual information loss suggested in [38] to further leverage overlapping similarities across the gathered biomedical domains. This way, we could correlate expert activation with certain domains without direct enforcement and concatenate the top-2 expert results at each layer. Additional local experiments showed that token-wise routing results in slightly higher-end performance, even on sentence-level tasks. Thus, we use the top-2 experts per token for our final five-domain model. Because our MoE extension can be costly in terms of VRAM, we also tried an MoE extension approach with a single transformer block in the middle instead of extending all 12 - hypothesizing that much of the multidomain benefit could be achieved for a small amount of extended MoE layers.
@@ 3 Results
The performance of our models on the validation and test datasets for CVD (Table 1) and COPD (Table 2) were summarized and compared to other leading sentence similarity models, as well as TF-IDF. We also evaluated the performance of GPT-3.5 on our generated validation and test datasets. GPT-3.5 represented an example of an LLM with a narrow representation of similar (Figure 2). GPT-3.5 regarded almost all examples that were input as dissimilar when asked whether the input examples were similar due to (correctly) identified minor differences between each input text. Given this, GPT-3.5 performance was excluded from model comparisons. We also do not show results from the MoE version of Mini as they did not score better than random on either validation dataset (0.67 $F1_{max}$).
@@ 4 Discussion
Our work advances the use of transformer language models by focusing on improving their domain-specific understanding and document-wide comprehension. We have shown that common pretrained models, including ChatGPT, cannot distinguish the differences in highly discriminative text inputs. Presumably, this phenotype in GPT-like LLMs is primarily caused by standard prompt construction and formatting. By including multiple distinct documents in a single prompt, the semantic similarity of similar tokens will prevent effective distinction between the documents, as portions of each document will attend highly to each other even if they are “different” as defined by a desired discrimination. With pretrained BERT-like model usage, presumably this poor performance comes from out-of-distribution tasks and a low ability to generalize with high discrimination. The sentence similarity approach offers a more effective summarization technique by inputting documents separately and allowing the model to construct descriptive numerical representations through contrastive objectives. As vector databases become increasingly prevalent for search and retrieval, the quality of these numerical representations becomes increasingly important. Our innovative framework, which incorporates contrastive learning, custom MNR and router losses, novel special tokens, MoE seeding, and extension, as well as various routing techniques, significantly enhances sentence similarity metrics compared to pretrained transformers. We leveraged co-citation networks to construct large datasets of similar abstracts and applied our framework to scientific literature, creating nuanced representations for discriminative subfields of biomedical domains.
@@
Without applying a threshold search for the F1 scores, all base pretrained models tested perform with random chance or worse when tasked with classifying papers as co-cited or not based on cosine similarity. By searching all possible cosine similarity thresholds for $F1_{max}$, pretrained models get a boost for fair comparison while creating a scenario that is unrealistic for actual inference, as a consistent similarity threshold is needed for actual use. Once compared fairly, our fine-tuned models performed better than their original counterparts. More specifically, our SciBERT SE and MoE variants performed equally on the CVD validation set with an $F1_{max}$ of 0.97, an average distance of 0.48, and an accuracy of 0.97. The Mini SE variant performed similarly with a $F1_{max}$ of 0.94, an average distance of 0.23, and an accuracy of 0.93. On the COPD validation set, Mini SE performed best with a $F1_{max}$ of 0.83, an average distance of 0.33, and an accuracy of 0.82. Following that is SciBERT SE with $F1_{max}$ 0.81 and SciBERT MoE with $F1_{max}$ 0.80. Both SciBERT variants resulted in an average distance of 0.49 and accuracies of 0.80 and 0.79, respectively. Importantly, our SciBERT variants consistently performed optimally with a high cosine similarity, implying that a standardized threshold near 0.98 or 0.99 could be utilized during inference. Further penalization terms on the average similarity between batches during training could enforce this threshold lower for different applications.
@@
Importantly, our MoE approach across all domains performs similarly to our individual-domain SE models. Additionally, the MoE seeding is fully scalable, appearing to enable $N$ experts and $N$ new special tokens given $N$ different datasets. This is further supported by our experiments shown in Table 3, where MoE models with five experts perform well on five domains, even with a single MoE layer. The substantial improvement in performance achieved by adding a single MoE layer highlights a remarkable benefit with little added computational cost. This is particularly promising for using our single-layer MoE extension with large pretrained models to train for multitask / multidomain tasks. Future experiments may find the optimal ratio of experts per domain alongside the correct discrimination of “domain” to create a one-size-fits-all vector embedding model at the scale of Semantic Scholar.
@@
Links to all training data, model weights, and code can be found at Github.
@@
The authors thank Katherine M. Nelson, Ph.D., for reviewing and commenting on drafts of the manuscript. This work was partly supported by grants from the University of Delaware Graduate College through the Unidel Distinguished Graduate Scholar Award (LH), the National Institutes of Health R01HL133163 (JPG), National Institutes of Health R01HL145147 (JPG), Indiana University (BBK), and NIH R01DK132090 (BBK).